If you randomly select a card from a well-shuffled standard deck of 52 cards, what is the probability that the card you select is a club or 7? (Your answer must be in the form of a reduced fraction.)


1
Expert's answer
2021-09-21T00:27:07-0400

There are 13 club in a deck, so the probability to choose some club is


P(club) = 13/52 = 1/4.


There are 4 7 in a deck, so the probability to choose a 7 is


P(7) = 4/52 = 1/13.


The probability to choose the 7 of club is


P(7 of club) = 1/52.


Therefore,


P(club or 7) = P(club) + P(7) - P(7 of club) = 1/4 + 1/13 - 1/52 = 4/13

=0.308.


So, the probability that the card you select is a club or 7 is about .308.